OpenStudy (anonymous):

Simplify -3 - [-2 + (5-6) -3] A. +3 B. -1 C. +1 D. -3

OpenStudy (anonymous):

@UditKulka @nincompoop @UsArmy3947 @BangkokGarrett

OpenStudy (anonymous):

-2+(-1)=-3 Can you explain this to me? I usually get confused with this and I want to get rid of this confusion. What if it were -2+1? How would the answer change in this equation.

OpenStudy (anonymous):

-3 - [-2 + (5 - 6) - 3] -3 - [-2 -1-3] -3 -[-6] -3 + 6 3 is your answer

OpenStudy (anonymous):

thank you UsArmy :)

OpenStudy (anonymous):

I still don't get it. Can either one of you two please explain? @UsArmy3947 @kelliegirl33

OpenStudy (anonymous):

-2 + (-1) = -3 a positive times a negative is a negative, therefore, this problem would be -2 - 1 = -3 -3 = -3 (correct) does that help ?

OpenStudy (anonymous):

@kelliegirl33 Thank you.

OpenStudy (anonymous):

do you understand, because I can explain it to you differently if you do not understand this way

OpenStudy (anonymous):

If you can, I will appreciate it :)

OpenStudy (anonymous):

-2 + (-1) = -3 This is basically about distributing through the parenthesis If there is a sign before the parenthesis and no number, it is understood that the unwritten number is 1. Let me show you... -2 +1(-1) = -3....basically this is the same as the above problem. now distribute the +1 through the parenthesis. +1(-1) = -1 (positive times a negative = a negative) I guess it is not much different then the other explanation...just remember if there is just a sign, the unwritten number is 1, and if it helps, put it in your problem.

OpenStudy (usarmy3947):

come on man now who took my medal

OpenStudy (usarmy3947):

u know im just helping to bring my score of problem solving up

OpenStudy (anonymous):

is that better :)

OpenStudy (usarmy3947):

yes thank you :)

OpenStudy (anonymous):

lol...no problem :)